Question: Evaluate: 1n(1/5sqrt e)
(A) 1/5
(B) -1/5
(C) -5
(D) 5
Thank you!

Answer by venugopalramana(3286)

Question: Evaluate: I SUPPOSE YOU MEAN.....PUT BRACKETS TO AVOID CONFUSION....1n((1)/(5sqrt e))=LN(1)-LN(5SQRT(E))=0-LN(5)-LN(E^0.5)
=-LN(5)-0.5LN(E)=-LN(5)-0.5...CHECK THE PROBLEM FOR PROPER TYPING
I THINK THE QUESTION IS LN(1/FIFTH ROOT OF E)
IT IS NOT SQUARE ROOT
IF NOTHING IS PUT IN V GROOVE OF THE SYMBOL IT IS SQUARE ROOT
IF 3 IS PUT IN V GROOVE OF THE SYMBOL IT IS CUBE ROOT
IF 4 IS PUT IN V GROOVE OF THE SYMBOL IT IS FOURTH ROOT
IF 5 IS PUT IN V GROOVE OF THE SYMBOL IT IS FIFTH ROOT...ETC....
I HOPE YOU KNOW 5 TH ROOT
IF 2*2*2*2*2=2^5=32
WE SAY 5TH.ROOT OF 32 =32^(1/5)= 32^0.2= 2...SO ASSUMING THAT WE GET
LN(1/FIFTH ROOT OF E)=LN(1/E^(1/5))=-LN(E^(1/5))=-(1/5)LN(E)=-1/5 IS THE ANSWER.
HENCE B IS THE ANSWER